OpenStudy (anonymous):

estimate the difference in energy between the 1st and 2nd bohr orbit for hydrogen atom.at what minimum atomic number would a transition from n=2 to n=1 energy level result in the emission of X-rays with ^=3.0*10^-8m?which hydrogen like species does thi atomic number corresponds to?

OpenStudy (anonymous):

@Frostbite

OpenStudy (anonymous):

@cambrige

OpenStudy (anonymous):

@Kayne @.Sam. @TheDoctorRules

OpenStudy (anonymous):

thnx buddy

OpenStudy (anonymous):

\[1/λ=1.09667∗10^{−7}(Z)(1/n _{i}^2−1/n _{f}^2)\]

OpenStudy (anonymous):

U got the answer from that??O_0

OpenStudy (anonymous):

doing it W8..

OpenStudy (anonymous):

Substitue your lamdba value and ni and nf values find the value of Z

OpenStudy (anonymous):

Wat do u get Z as?? Take ur time!1

OpenStudy (anonymous):

This equation was for the 2nd part of the problem!!

OpenStudy (anonymous):

For the first part use

OpenStudy (anonymous):

\[\Delta E={R _{H}}({1/n1^2-1/n2^2})\]

OpenStudy (anonymous):

i got that Z=0.101

OpenStudy (anonymous):

Where \[R _{H}=-2.18*10^{-18} J\]

OpenStudy (anonymous):

Sure??

OpenStudy (anonymous):

Is 3*10^-8 the wavelength or frequency??

OpenStudy (anonymous):

it,s wavelength

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

I did a mistake...Its 1.09*10^7 not 10^-7

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

Okk then i got Z=1.01*10^-2

OpenStudy (anonymous):

U get Z=4 so It is Helium!!

OpenStudy (anonymous):

U dint multiply 1.01 with 4!!

OpenStudy (anonymous):

And how did u get 10^-2???O_O

OpenStudy (anonymous):

\[10^8/3=1.09667*10^7*Z*(1/n _{i}^2-1/n _{f}^2)\]

OpenStudy (anonymous):

Solve this and get me the ans!!

OpenStudy (anonymous):

Its square of n(i) and n(f) giving u 4 in the denominator!!lol

OpenStudy (anonymous):

Using xact vlues u get approx 4.0528 i.e 4! So this atomic number belongs to helium!! Get me ??

OpenStudy (anonymous):

@hamza10 u there??

OpenStudy (anonymous):

@cambrige sorry i was a little busy

OpenStudy (anonymous):

yes i got u